OpenStudy (anonymous):

Solve each equation by multiplying. z/8 = 12

jimthompson5910 (jim_thompson5910):

Multiply both sides by 8 to isolate z z/8 = 12 8*(z/8) = 12*8 z = 96

OpenStudy (anonymous):

z/8(8)=12(8) z=96